OpenStudy (gabylovesyou):

Sergio has two similar cylindrical pots, Pot A and Pot B. The radius of Pot A is 15 inches, and the radius of Pot B is 3 inches. What is the ratio of the volume of Pot A to the volume of Pot B? 10:1 25:1 125:1 5:1

OpenStudy (quietus):

25:1, if assuming that the height is the same.

OpenStudy (goformit100):

We know the Volume of Cylinder is 

OpenStudy (quietus):

volume of cylinder = \[\pi \times r^2 \times height\]

OpenStudy (anonymous):

geez everyone so eager to help her lol

OpenStudy (gabylovesyou):

@$w3G_Godd im special ;D

OpenStudy (goformit100):

For Cylinder pot A, we will have  And for Second it will be